# SHAP values via sparse Fourier representation

### v2

@@ Abstract
SHAP (SHapley Additive exPlanations) values are a widely used method for local feature attribution in interpretable and explainable AI. We propose an efficient two-stage algorithm for computing SHAP values in both black-box setting and tree-based models. Motivated by spectral bias in real-world predictors, we first approximate models using compact Fourier representations, exactly for trees and approximately for black-box models. In the second stage, we introduce a closed-form formula for exactly computing SHAP values using the Fourier representation, that “linearizes” the computation into a simple summation and is amenable to parallelization. As the Fourier approximation is computed only once, our method enables amortized SHAP value computation, achieving significant speedups over existing methods and a tunable trade-off between efficiency and precision.
@@ 1 Introduction
A widely used method for deriving attributions (importances) is the notion of SHapley Additive exPlanations, commonly referred to simply as SHAP values. Originally, the notion of Shapley values was introduced in the seminal work of Shapley [1952] in the context of cooperative game theory. The Shapley value is a mathematically well-founded and “fair” way of distributing a reward among all the members of a group playing a cooperative game and it is computed based on the rewards that would be received for all possible coalitions. The Shapley value is the unique way of distributing the reward that satisfies several reasonable mathematical properties that capture a notion of fairness [Shapley, 1952]. In the context of machine learning and statistics, the players become features, the reward is the prediction of the predictor $h$ and the SHAP value is the “contribution” or “influence” of that feature on the prediction. Shapley values are widely used due to their mathematical soundness and desirable properties Gromping [2007], Štrumbelj et al. [2009], Owen [2014], Datta et al. [2016], Owen and Prieur [2017], Lundberg and Lee [2017], Lundberg et al. [2020] Aas et al. [2021].
Despite their prevalence, computing SHAP values is challenging, as it involves an exponential summation, i.e. a summation over exponentially many terms , see Equation 1. This is because the formula accounts for a feature’s importance in the context of all possible “coalitions” of other features, and therefore the formula covers all possible subsets of other features. Therefore, approximating them and speeding up the computation has received attention in a variety of settings. SHAP value computation can easily dominate the computation time of industry-level machine learning solutions on datasets with millions or more entries [Yang, 2021]. Yang [2021] point out that industrial applications sometimes require hundreds of millions of samples to be explained. Examples include feed ranking, ads targeting, and subscription propensity models. In these modeling pipelines, spending tens of hours in model interpretation becomes a significant bottleneck [Yang, 2021] and one usually needs to resort to multiple cores and parallel computing.
@@
As opposed to the tree setting, which is a “white box” setting, in the model-agnostic a.k.a black-box setting, we only have query access to the model. Here, our only means of access to the predictor is that we can pick an arbitrary $x\in\mathcal{X}^{n}$ and query the predictor for its value $h(x)$. The usual approach here is to approximate the exponential sum of the SHAP value computations using stochastic sampling [Covert and Lee, 2020, Mitchell et al., 2022a, Lundberg and Lee, 2017]. In this setting, Covert and Lee [2020], Mitchell et al. [2022a] provide sampling methods that require fewer queries to the black-box compared to vanilla KernelShap [Lundberg and Lee, 2017] for equal approximation accuracy. FastShap Jethani et al. [2021], which introduces a method for estimating Shapley values in a single forward pass using an end-to-end learned explainer neural network model. Our algorithm FourierShap falls into the query-access black-box setting as well.
However, we take a different approach. We are guided by the key insight that many models used in practice have a “spectral bias”. Yang and Salman [2019], Valle-Perez et al. [2018] provably and experimentally show that fully connected neural networks with binary (zero-one) inputs learn low-degree – and therefore sparse – functions in a basis called the Walsh-Hadamard a.k.a Fourier basis. It is well known that the Walsh-Hadamard transform (WHT) of an ensemble of $T$ trees of depth $d$ is also of degree at most $d$ and moreover, $k=O(T4^{d})$-sparse [Kushilevitz and Mansour, 1993, Mansour, 1994].
@@
Our contributions: Guided by the aforementioned insights, we provide an algorithm to compute SHAP values, using the Fourier representation of the tree or black-box model. We first approximate the black-box function by taking its sparse Fourier transform. We theoretically justify, and show through extensive experiments, that for many real-world models such as fully connected neural networks and (ensembles of) trees this representation is accurate. Subsequently, we prove that SHAP values for a single Fourier basis function admit a closed-form expression not involving an exponential summation. Therefore, using the Fourier representation we overcome the exponential sum and can utilize compute power effectively to compute SHAP values. Furthermore, the closed-form expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. The Fourier approximation step is only done once, therefore FourierShap amortizes the cost of computing explanations for many inputs. Subsequently, SHAP computations using the Fourier approximation are orders of magnitude faster compared to other black-box SHAP approximation methods such as KernelShap and other variations of it which all involve a computationally expensive optimization. We show speedups over other methods such as DeepLift [Shrikumar et al., 2017] and FastShap [Jethani et al., 2021] as well. In addition to the speedup, our algorithm enables a reliable continuous trade-off between computation and accuracy, controlled in a fine manner by the sparsity of the Fourier approximation.
@@ 2.1 Shapley values
In game theory, a cooperative game is a function $v:2^{[n]}\rightarrow\mathbb{R}$ that maps a subset (coalition) $S\subseteq[n]$ of a group of players $[n]=\{1,\dots,n\}$ to their total reward (when they cooperate). If all the players cooperate they win a total reward of value $v([n])$. The main question is how they would distribute this reward among themselves. Shapley [1952] resolved this problem by a deriving a unique value based on “fairness axioms” proposed in his seminal work [Shapley, 1952]. The Shapley value of player $i\in[n]$ is:
@@
Intuitively, one can view the term $v(S\cup\{i\})-v(S)$ as the marginal contribution of player $i$ when they are added to the coalition $S$. This marginal value is weighted according to the number of permutations the leading $|S|$ players and trailing $n-|S|-1$ players can form.
@@
Since we will be using the well-known KernelShap [Lundberg and Lee, 2017] and its variant LinRegShap [Covert and Lee, 2020] as a baseline we briefly review their method here. Lundberg and Lee [2017] propose the “least squares characterization” of SHAP values. They prove that SHAP values are the solution to the following minimization problem:
@@
Later, we also compare our method to FastSHAP [Jethani et al., 2021], a model-agnostic algorithm for computing SHAP values. In FastSHAP, a parametric explainer $\phi$ (e.g., MLP) is trained to directly generate SHAP values given data samples. Since training is only done once, this, similar to us, amortizes the cost of generating SHAP values across multiple instances. Computing SHAP values only requires a forward pass on the trained model and therefore can be done very quickly.
@@ 2.2 Fourier representations
Here we review the notions of the Fourier basis and what we mean by sparsity. We will later use sparse Fourier representation of functions to compute SHAP values. In this work, we focus on the setting where the inputs to the black-box function (predictor) are binary, i.e., $\mathcal{X}^{n}=\{0,1\}^{n}$. This means, we assume we have binary features and/or categorical features, through standard one-hot representations  [^1] Continuous features can be discretized into quantiles, enabling their transformation into categorical features.. The Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows: $h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$.
@@ 4 Computing SHAP values with Fourier representation of functions
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be some predictor that we assume to be $k$-sparse. Assume $[n]\triangleq\{1,\dots,n\}$ as the set of features and let $x^{*}\in\{0,1\}^{n}$ be the instance we are explaining. Aligned with [Lundberg and Lee, 2017], we use $v_{h}(S)\triangleq\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}h(x^{*}_{S},x_{[n]\setminus S})$ as the value function, which is the average prediction when fixing $x_{S}^{*}$ in the (background) dataset $\mathcal{D}$. We compute the Shapley values for a single Fourier basis function $\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$ according to Equation 1:
@@
Theorem 2 gives us a computationally efficient way to go from a Fourier representation of a function $h$ to SHAP values. The SHAP values from this equation are exact, i.e., as long as the Fourier representation is exact, the SHAP values are also precise values. This is in contrast to KernelSHAP, where the SHAP values are approximated using stochastic sampling and one needs to check for convergence to ensure the accuracy. The approximation in our method is constrained to computing the sparse Fourier representation, only in the black-box setting.
Most importantly, the sum in Equation 5 is tractable. This is because we overcome the exponential sum involved in Equation 1 by analytically computing the sum, with a combinatorial argument, for a single Fourier basis function. We show the number of flops that are required to compute SHAP values in Theorem 2 to be asymptotically equal to $\Theta(n\cdot|\mathcal{D}|\cdot k)$. We note that the $|\mathcal{D}|$ and $k$ factors in the asymptotic computational complexity arise from the two summations present in Equation 5. Through this expression, we are able to “linearize” the computation of SHAP values to a summation over the Fourier coefficients and background dataset. This allows us to maximally utilize the parallelization on multiple cores and/or GPUs to speed up the computation significantly. Therefore, in the presence of multiple cores or a GPU, we can get a speedup equal to the level of parallelization, as each core or worker can compute one part of this summation.
We implement our algorithm called FourierShap using JAX [Bradbury et al., 2018], which allows for fast vectorized operations on GPUs. Each term inside the summations of Equation 5 can be implemented with simple vector operations. Furthermore, summations over the $k$ different frequencies in the support of $h$ and also background data points can both be efficiently implemented and parallelized using this library using its vmap operator. We perform all upcoming experiments on a single GPU. Nevertheless, we believe faster computation can also be achieved by crafting dedicated code designed to efficiently compute Equation 5 on GPU.
@@ 5 Experiments
Finally, we compare against FastShap [Jethani et al., 2021], which is the closest to our setting, i.e., the model agnostic black-box, amortized setting. We provide higher accuracy for similar values of speedup compared to FastShap in 3/4 settings. More importantly though, we can see by controlling the sparsity parameter $k$ of the Fourier function approximation, we can control the tradeoff between accuracy and speed in a reliable and fine-grained manner. Whereas for FastShap the accuracy of the SHAP values relies on the functional approximation properties of the MLP which directly produces SHAP values. As seen in Figure 2, the MLP can behave in unpredictable ways. In this figure we can see that increasing the depth of the model (and hence the approximation capability of the MLP) does not have a reliable effect on SHAP accuracy. Finally, in FastShap for a different choice of background dataset a new MLP model has to be trained from scratch. Since, the MLP approximates the process of directly computing SHAP values. Whereas we can support different sets of background datasets since our Fourier functional approximation is done on the black box predictor, and not the whole process of producing SHAP values in an end-to-end manner (FastShap).
@@
To demonstrate our method’s ability in fast computation of SHAP values in this setting, we compute SHAP values for random forests fitted on all aforementioned real-world datasets. We compare FourierSHAP’s speedup over TreeSHAP [Lundberg et al., 2020], to FastTreeSHAP [Yang, 2021], a fast implementation of TreeSHAP, the GPU implementation of TreeSHAP [Mitchell et al., 2022b] and PLTreeSHAP [Zern et al., 2023]. To the best of our knowledge, these are the fastest available frameworks for computation of the “interventional” SHAP values. Figure 3 shows the superior speed of our method over all state-of-the-art algorithms in most settings, which weakens by the increment of depth and number of features, resulting in more frequencies and computation overhead. Note that the SHAP values computed by all methods are precise and identical to the values produced by TreeSHAP, which is to be expected for FourierSHAP, given access to exact Fourier representation of random forests.
@@ Conclusions
We illustrated in theory and practice how many black-box functions can be represented or efficiently approximated by a compact Fourier representation. We proved that SHAP values of Fourier basis functions admit a closed form expression, and therefore, we can compute SHAP values efficiently using the compact Fourier representation. Moreover, this closed form expression is amenable to parallelization. These two factors helped us in gaining speedups of 10-10000x over baseline methods for the computation of SHAP values.

### v3

@@ Abstract
SHAP (SHapley Additive exPlanations) values are a widely used method for local feature attribution in interpretable and explainable AI. We propose an efficient two-stage algorithm for computing SHAP values in both black-box setting and tree-based models. We assume the black-box predictor or tree model accepts binary (zero-one) features. Motivated by spectral bias in real-world predictors, we first approximate the predictor using compact Fourier representations, exactly for trees and approximately for black-box models. In the second stage, we introduce a closed-form formula for exactly computing SHAP values using the Fourier representation, that “linearizes” the computation into a simple summation and is amenable to parallelization. As the Fourier approximation is computed only once, our method enables amortized SHAP value computation, achieving significant speedups over existing methods and a tunable trade-off between efficiency and precision.
@@ 1 Introduction
A widely used method for deriving attributions (importances) is the notion of SHapley Additive exPlanations, commonly referred to simply as SHAP values. Originally, the notion of Shapley values was introduced in the seminal work of Shapley [1952] in the context of cooperative game theory. The Shapley value is a mathematically well-founded and “fair” way of distributing a reward among all the members of a group playing a cooperative game and it is computed based on the rewards that would be received for all possible coalitions. The Shapley value is the unique way of distributing the reward that satisfies several reasonable mathematical properties that capture a notion of fairness [Shapley, 1952]. In the context of machine learning and statistics, the players become features, the reward is the prediction of the predictor $h$ and the SHAP value is the “contribution” or “influence” of that feature on the prediction. Shapley values are widely used due to their mathematical soundness and desirable properties (Gromping [2007], Štrumbelj et al. [2009], Owen [2014], Datta et al. [2016], Owen and Prieur [2017], Lundberg and Lee [2017], Lundberg et al. [2020] Aas et al. [2021]).
Despite their prevalence, computing SHAP values is challenging, as it involves an exponential summation, i.e., a summation over exponentially many terms , see Equation 1. This is because the formula accounts for a feature’s importance in the context of all possible “coalitions” of other features, and therefore the formula covers all possible subsets of other features. Therefore, approximating them and speeding up the computation has received attention in a variety of settings. SHAP value computation can easily dominate the computation time of industry-level machine learning solutions on datasets with millions or more entries [Yang, 2021]. Yang [2021] point out that industrial applications sometimes require hundreds of millions of samples to be explained. Examples include feed ranking, ads targeting, and subscription propensity models. In these modeling pipelines, spending tens of hours in model interpretation becomes a significant bottleneck [Yang, 2021] and one usually needs to resort to multiple cores and parallel computing.
@@
As opposed to the tree setting, which is a “white box” setting, in the model-agnostic a.k.a black-box setting, we only have query access to the model. Here, our only means of access to the predictor is that we can pick an arbitrary $x\in\mathcal{X}^{n}$ and query the predictor for its value $h(x)$. The usual approach here is to approximate the exponential sum of the SHAP value computations using stochastic sampling [Covert and Lee, 2020, Mitchell et al., 2022a, Lundberg and Lee, 2017]. In this setting, Covert and Lee [2020], Mitchell et al. [2022a] provide sampling methods that require fewer queries to the black-box compared to vanilla KernelShap [Lundberg and Lee, 2017] for equal approximation accuracy. FastShap Jethani et al. [2021], which introduces a method for estimating Shapley values in a single forward pass using an end-to-end learned explainer neural network model. Our algorithm FourierShap falls into the query-access black-box setting as well. However, we take a different approach. We are guided by the key insight that many models used in practice have a “spectral bias”. Yang and Salman [2019], Valle-Perez et al. [2018] provably and experimentally show that fully connected neural networks with binary (zero-one) inputs learn low-degree – and therefore sparse – functions in a basis called the Walsh-Hadamard a.k.a Fourier basis. It is well known that the Walsh-Hadamard transform (WHT) of an ensemble of $T$ trees of depth $d$ is also of degree at most $d$ and moreover, $k=O(T4^{d})$-sparse [Kushilevitz and Mansour, 1993, Mansour, 1994].
@@
Our contributions: Guided by the aforementioned insights, we provide an algorithm to compute SHAP values using the Fourier representation of the tree or black-box model. We first approximate the black-box function by taking its sparse Fourier transform. We theoretically justify, and show through extensive experiments, that for many real-world models such as fully connected neural networks and (ensembles of) trees this representation is accurate. Subsequently, we prove that SHAP values for a single Fourier basis function admit a closed-form expression not involving an exponential summation. Therefore, using the Fourier representation we overcome the exponential sum and can utilize compute power effectively to compute SHAP values. Furthermore, the closed-form expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. The Fourier approximation step is only done once, therefore FourierShap amortizes the cost of computing explanations for many inputs. Subsequently, SHAP computations using the Fourier approximation are orders of magnitude faster compared to other black-box SHAP approximation methods such as KernelShap and other variations of it which all involve a computationally expensive optimization. We show speedups over other methods such as DeepLift [Shrikumar et al., 2017] and FastShap [Jethani et al., 2021] as well. In addition to the speedup, our algorithm enables a reliable continuous trade-off between computation and accuracy, controlled in a fine manner by the sparsity of the Fourier approximation.
@@ 2.1 Shapley values
In game theory, a cooperative game is a function $v:2^{[n]}\rightarrow\mathbb{R}$ that maps a subset (coalition) $S\subseteq[n]$ of a group of players $n]=\{1,\dots,n\}$ to their total reward (when they cooperate). If all the players cooperate they win a total reward of value $v([n])$. The main question is how they would distribute this reward among themselves. Shapley [1952] resolved this problem by a deriving a unique value based on “fairness axioms” proposed in his seminal work [Shapley, 1952]. The Shapley value of player $i\in[n]$ is:
@@
Intuitively, one can view the term $v(S\cup\{i\})-v(S)$ as the marginal contribution of player $i$ when they are added to the coalition $S$. This marginal value is weighted according to the number of permutations the leading $|S|$ players and trailing n|S|1n-|S|-1 players can form.
@@
Since we will be using the well-known KernelShap [Lundberg and Lee, 2017] and its variant LinRegShap [Covert and Lee, 2020] as a baseline, we briefly review their method here. Lundberg and Lee [2017] propose the “least squares characterization” of SHAP values. They prove that SHAP values are the solution to the following minimization problem:
@@
Later, we also compare our method to FastSHAP [Jethani et al., 2021], a model-agnostic algorithm for computing SHAP values. In FastSHAP, a parametric explainer $\phi$ (e.g., MLP) is trained to directly generate SHAP values given data samples. Since training is only done once, this, similar to us, it amortizes the cost of generating SHAP values across multiple instances. Computing SHAP values only requires a forward pass on the trained model and therefore can be done very quickly.
@@ 2.2 Fourier representations
Here we review the notions of the Fourier basis and what we mean by sparsity. We will later use sparse Fourier representation of functions to compute SHAP values. In this work, we focus on the setting where the inputs to the black-box function (predictor) are binary, i.e., $\mathcal{X}^{n}=\{0,1\}^{n}$. This means, we assume we have binary features and/or categorical features, through standard one-hot representations Continuous features can be discretized into quantiles, enabling their transformation into categorical features. The Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows: $h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$.
@@ 4 Computing SHAP values with Fourier representation of functions
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be some predictor that we assume to be $k$-sparse. Assume $n]\triangleq\{1,\dots,n\}$ as the set of features and let $x^{*}\in\{0,1\}^{n}$ be the instance we are explaining. Aligned with [Lundberg and Lee, 2017], we use $v_{h}(S)\triangleq\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}h(x^{*}_{S},x_{[n]\setminus S})$ as the value function, which is the average prediction when fixing $x_{S}^{*}$ in the (background) dataset $\mathcal{D}$. We compute the Shapley values for a single Fourier basis function $\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$ according to Equation 1:
@@
Intuitively, feature $i$ contributes to the prediction only if the Fourier basis function $\psi_{f}$ is sensitive to the value of feature $i$ i.e. $f_{i}=1$ and otherwise its contribution is zero. The SHAP values is also dependent on the parity of the number of differences between the query and background samples within support of $f$: many positive and negative terms cancel during the exponential summation. The key combinatorial insight is that we can count the number of cancellations in the exponential sum without computing it and therefore this collapses into a closed-form factor depending on the parity of the “overlap set” $A$.
@@
Theorem 2 provides a computationally efficient method for deriving SHAP values from a Fourier representation of a function $h$. The resulting SHAP values are exact, that is, if the Fourier representation is precise, then the SHAP values are also exact. This contrasts with KernelSHAP, which approximates SHAP values via stochastic sampling, requiring convergence checks to ensure accuracy. In our method, any approximation arises solely from estimating a sparse Fourier representation, and only in the black-box setting.
More importantly, the sum in Equation 5 is tractable. This is because we overcome the exponential sum involved in Equation 1 by analytically computing the sum, with a combinatorial argument, for a single Fourier basis function. We show that the number of flops that are required to compute SHAP values in Theorem 2 is asymptotically equal to $\Theta(n\cdot|\mathcal{D}|\cdot k)$. We note that the $|\mathcal{D}|$ and $k$ factors in the asymptotic computational complexity arise from the two summations present in Equation 5. Through this expression, we are able to “linearize” the computation of SHAP values to a summation over the Fourier coefficients and the background data set. This allows us to maximize parallelization on multiple cores and/or GPUs to speed up the computation significantly. Therefore, in the presence of multiple cores or a GPU, we can get a speedup equal to the level of parallelization, as each core or worker can compute one part of this summation.
We implement our algorithm called FourierShap using JAX [Bradbury et al., 2018], which allows fast vectorized operations on GPUs. Each term within the summations of Equation 5 can be implemented with simple vector operations. Furthermore, summations over the $k$ different frequencies in the support of $h$ and also background data points can both be efficiently implemented and parallelized using this library using its vmap operator. We perform all upcoming experiments on a single GPU. Nevertheless, we believe faster computation can also be achieved by crafting dedicated code designed to efficiently compute Equation 5 on GPU.
@@ 5 Experiments
Finally, we compare against FastShap [Jethani et al., 2021], which is the most comparable to our setting, i.e., model-agnostic with amortized cost. We provide a higher accuracy for similar speedup values compared to FastShap in 3/4 settings. More importantly, we can see that by controlling the sparsity parameter $k$ of the Fourier function approximation, we can control the tradeoff between accuracy and speed in a reliable and fine-grained manner. Whereas for FastShap, the accuracy of the SHAP values relies on the functional approximation properties of the MLP which directly produce SHAP values. As seen in Figure 2, the MLP can behave in unpredictable ways. In this figure, we can see that increasing the depth of the model (and hence the approximation capability of the MLP) does not have a reliable effect on SHAP accuracy. Finally, in FastShap, for a different choice of background dataset, a new MLP model has to be trained from scratch since the MLP approximates the process of directly computing SHAP values. In contrast, we can support different sets of background datasets since our Fourier functional approximation is performed on the black-box predictor, and not on the whole process of producing SHAP values end-to-end (FastShap).
@@
To demonstrate our method’s ability in fast computation of SHAP values in this setting, we compute SHAP values for random forests fitted on all aforementioned real-world datasets. We compare FourierSHAP’s speedup over TreeSHAP [Lundberg et al., 2020], to FastTreeSHAP [Yang, 2021], a fast implementation of TreeSHAP, the GPU implementation of TreeSHAP [Mitchell et al., 2022b] and PLTreeSHAP [Zern et al., 2023]. To the best of our knowledge, these are the fastest available frameworks for computation of the “interventional” SHAP values. Figure 3 shows the superior speed of our method over all state-of-the-art algorithms in most settings, which weakens with increased depth and number of features, resulting in a higher count of frequencies and greater computational overhead. Note that the SHAP values computed by all methods are precise and identical to the values produced by TreeSHAP, which is to be expected for FourierSHAP, given access to exact Fourier representation of random forests.
@@ Conclusions
We illustrated in theory and practice how many black-box functions can be represented or efficiently approximated by a compact Fourier representation. We proved that SHAP values of Fourier basis functions admit a closed form expression, and therefore, we can compute SHAP values efficiently using the compact Fourier representation. Moreover, this closed form expression is amenable to parallelization. These two factors helped us in gaining speedups of 10-10000x over baseline methods for the computation of SHAP values. We discuss the limitations & future works in Appendix F.
Acknowledgments
This work was supported by the Swiss National Science Foundation (SNSF) through National Centre of Competence in Research (NCCR) Catalysis (Grant No. 180544). We thank Bhavya Sukhija for valuable feedback and help during the review and rebuttal process.